# Targeted Attack Improves Protection against Unauthorized Diffusion Customization

### v4

@@ Abstract
Diffusion models build a new milestone for image generation yet raising public concerns, for they can be fine-tuned on unauthorized images for customization. Protection based on adversarial attacks rises to encounter this unauthorized diffusion customization, by adding protective watermarks to images and poisoning diffusion models. However, current protection, leveraging untargeted attacks, does not appear to be effective enough. In this paper, we propose a simple yet effective improvement for the protection against unauthorized diffusion customization by introducing targeted attacks. We show that by carefully selecting the target, targeted attacks significantly outperform untargeted attacks in poisoning diffusion models and degrading the customization image quality. Extensive experiments validate the superiority of our method on two mainstream customization methods of diffusion models, compared to existing protections. To explain the surprising success of targeted attacks, we delve into the mechanism of attack-based protections and propose a hypothesis based on our observation, which enhances the comprehension of attack-based protections. To the best of our knowledge, we are the first to both reveal the vulnerability of diffusion models to targeted attacks and leverage targeted attacks to enhance protection against unauthorized diffusion customization. Our code is available on GitHub: https://github.com/psyker-team/mist-v2.
@@ 2 Background
As in existing works, we only consider LDM because it is the basis for most mainstream (unauthorized) diffusion customization. Notably, there are various architecture variants for diffusion models nowadays, including IF and DALLE. However, neither of these architectures supports customizations, thus avoiding concerns of unauthorized customization. Hence, in the rest of this paper we focus on LDM and its customization.
Diffusion Customization Diffusion customization (or diffusion personalization (Zhang et al., 2024)) refers to customizing the generated images by diffusion models according to the references from user inputs. These user inputs cover reference images (Meng et al., 2021), text description (Saharia et al., 2022), sketches (Zhang & Agrawala, 2023), styles (Ruiz et al., 2023; Zhang et al., 2023), and identities (Ruiz et al., 2023; Ye et al., 2023; Wang et al., 2024). Among these customization methods, LoRA+DreamBooth (Hu et al., 2021; Ruiz et al., 2023) and SDEdit (Meng et al., 2021) are two mainstream methods that cover most of the customization practice (See our brief survey in Appendix D). Hence, we follow the setup of existing works (Van Le et al., 2023; Liang et al., 2023; Xue et al., 2023) to focus on these two methods.
Adversarial Attacks & Protections against Unauthorized Diffusion Customization Adversarial attacks on Diffusion Models (Salman et al., 2023; Liang et al., 2023; Shan et al., 2023a; Liang & Wu, 2023; Van Le et al., 2023; Zhao et al., 2023; Xue et al., 2023) have been used to protect private images against unauthorized diffusion customization. Specifically, these attacks add tiny adversarial perturbations to private images as protections. Diffusion models customized (fine-tuned (Hu et al., 2021; Ruiz et al., 2023) or referring to (Meng et al., 2021)) the protected images will produce images of degraded quality in customized generation. This will penalize the attempt to customize diffusion models on unauthorized private images.
It is noticeable that all existing protections are based on untargeted attacks on diffusion models. Those untargeted objectives maximize the score-function error in Equation 2 by optimizing the image perturbation $\eta$ within the perturbation budget $\zeta$, which is formulated by Equation 3.
@@ 3.1 Attacking with Consistent Errors
$&\max_{\eta}\|p_{\theta}(y|x^{\prime})-\bm{e}_{true}\|_{2}^{2},\quad\quad\quad\quad&\text{(untargeted attack)}\\ &{\color[rgb]{1,.5,0}\min_{\eta}}\|p_{\theta}(y|x^{\prime})-{\color[rgb]{1,.5,0}\bm{e}_{target}}\|_{2}^{2},\quad\quad\quad\quad&\text{(targeted attack)}\end{aligned}\right.$
@@
$&\max_{\eta}\mathbb{E}_{t}\mathbb{E}_{z^{\prime}(t)|z^{\prime}(0)}\|s_{\theta}(z^{\prime}(t),t)-\nabla_{z^{\prime}(t)}\log p_{t}(z^{\prime}(t)|z^{\prime}(0))\|^{2}_{2},\quad\quad\quad\quad&\text{(untargeted attack)}\\ &{\color[rgb]{1,.5,0}\min_{\eta}}\mathbb{E}_{t}\mathbb{E}_{z^{\prime}(t)|z^{\prime}(0)}\|s_{\theta}(z^{\prime}(t),t)-{\color[rgb]{1,.5,0}\mathcal{T}}\|^{2}_{2},\quad\quad\quad\quad&\text{(targeted attack)}\end{aligned}\right.$
@@ 4 Experiments
We compare ACE/ACE+/ACE to existing protections against unauthorized diffusion customization in two main diffusion customization methods, LoRA+DreamBooth (Hu et al., 2021; Ruiz et al., 2023) and SDEdit (Meng et al., 2021). We also evaluate our protection on DreamBooth, detailed in Appendix B.3. Our methods outperform baselines both in quantitative metrics and qualitative visualization (Section 4.2). We also conduct user studies in an artist community, which validates the superiority of our protections (Section 4.3). Furthermore, we test the transferability (Section 4.4) and robustness to purification (Section 4.5) of ACE/ACE+. We also investigate the computational cost of our methods and baselines (Appendix B.1). Our ablation studies focus on the effect of the objective function, target selection (Appendix B.4.1), different text prompts (Appendix B.4.2) and perturbation magnitudes (Appendix B.4.3) to the protection effectiveness.
@@ 4.1 Experimental Setups
Baselines We compare ACE/ACE+ with existing methods, including AdvDM (Liang et al., 2023), Encoder Attack and Diffusion Attack in PhotoGuard (Salman et al., 2023), and ASPL (Anti-DreamBooth) (Van Le et al., 2023). Encoder Attack and Diffusion Attack in PhotoGuard (Salman et al., 2023) are denoted by PG and PG+, respectively.
@@ 4.2 Overall Result
In addition to quantitative evaluation, we visualize the customization output images under different protection. Figure 1 showcases output images of LoRA and SDEdit under the protection of baselines and ACE/ACE+. We visualize the output images of ACE as an ablation study in Figure 10. Notably, all baseline methods only add chaotic patterns to some parts of the output image. Diffusion customization still succeeds in generating images with correct contents and styles. In contrast, ACE/ACE+ covers the whole output image with high contrast and overly busy patterns, making the image completely unusable. This advantage distinguishes ACE/ACE+ from existing protection against unauthorized diffusion customization. More visualization is given in Appendix E.
@@ 4.3 User Study
Protections against unauthorized diffusion customization aim at making the output image of unauthorized customization unusable in graphic art industries. The gold standard for evaluating protection is user studies by graphic artists. Hence, we conduct a survey in an artist community on the comparative effectiveness of two most competitive protections: ASPL (Van Le et al., 2023), and ACE. Basically, our survey asks participants to pick the image with the worst quality out of two customization output images under the protection of ASPL, and ACE, respectively. The strongest protection should make the output image of the worst quality. We omit the details to Appendix D. The survey covers 1304 artists, whose result shows that ACE produces the worst customization output image among $55\%$ of the examples.
@@ 4.4 Transferability
Protections against unauthorized diffusion customization are based on white-box and model-dependent adversarial attacks. Hence, it is important to investigate the transferability of ACE/ACE+ over different diffusion models to validate their utility. We follow existing literature (Liang et al., 2023; Xue et al., 2023) to evaluate this transferability on three text-guided diffusion models: Stable Diffusion 1.4, Stable Diffusion 1.5, and Stable Diffusion 2.1. As mentioned in Section 2 and Section 4.1 and shown in Appendix D, we only consider these SD-family models because they cover most of unauthorized diffusion customization.
In this experiment, backbone means the model used to generate adversarial examples and victim means the model used in few-shot generation. We pick 100 images from CelebA-HQ, 20 in each of 5 groups, to generate adversarial examples on three backbone models and run SDEdit and LoRA with these adversarial examples on three victim models. The strength of SDEdit is set to 0.4 for visualizing the differences. Other experimental setups stay the same with Section 4.1. Table 2 shows that ACE/ACE+ can degrade diffusion customization even when the victim is different from the backbone. We omit the detailed explanation of the result to Appendix E.2.
@@ 5 Discussion: How can targeted attack beats untargeted attack?
While ACE/ACE+/ACE beat baselines in protecting private images from unauthorized diffusion customization empirically, it is non-trivial to explain our success. In this section, we discuss the process of fine-tuning diffusion models on protected images and provide one view of analysis on how these attack-based protections work and why ACE/ACE+/ACE outperform existing methods.
@@
2) $\mathcal{B}_{spl}$ is the score-function error of clean image $x$ in the original diffusion model $\theta^{\prime}$.
[…]
@@ Appendix A Implementation and Experimental Details
[…]
@@ B.2 Robustness to purification methods
Qualitatively, Table 3 also visualizes the output images of SDEdit and LoRA referring to the protected images processed by different purification methods. For both method, they still have strong performance under most of the cases except for Gaussian Noise($\sigma=8$) and JPEG compression (quality$=20$). However, in the exception cases, the defense has added visible degradation to the image, which also heavily affect both LoRA and SDEdit process. For example, LoRA learns to produce images comprised of small squares due to a hard compression of quality 20. And SDEdit produces images of visible Gaussian noise when adding Gaussian noise of $\sigma=8$. It’s noteworthy that both ACE and ACE+ seems to be strengthened after SR purification, which is an intriguing phenomena.
@@ B.3 Performance on Dreambooth
B.4 Ablation Studies
B.4.1 Targets and Objectives
Connections between ACE/ACE+ and other targeted attacks ACE/ACE+ are not the first targeted attack on LDM. ASPL-T is a general targeted form of attacks (Van Le et al., 2023). Here, we compare the objective function of ACE and ASPL-T  [^13] The original paper does not provide the concrete objective function of ASPL-T so we refer to its implementation on GitHub: https://github.com/VinAIResearch/Anti-DreamBooth/blob/main/attacks/aspl.py:
Effects of Targets and Objectives We conduct ablation studies on the effect of targets and objectives. For targets, we use two natural images, one human-face photo from CelebA-HQ and one artwork from WikiArt, as alternative images for the target $\bm{\mathcal{T}}$. This is recommended by Van Le et al. (2023). For objectives, we use ASPL-T as the alternative objective. The experiment setup follows the LoRA setup in Section 4.1.
@@ Appendix C Limitations & Societal Impacts
Appendix D Toy Survey on Diffusion Customization

### v5

@@ Abstract
Diffusion models build a new milestone for image generation yet raising public concerns, for they can be fine-tuned on unauthorized images for customization. Protection based on adversarial attacks rises to encounter this unauthorized diffusion customization, by adding protective watermarks to images and poisoning diffusion models. However, current protection, leveraging untargeted attacks, does not appear to be effective enough. In this paper, we propose a simple yet effective improvement for the protection against unauthorized diffusion customization by introducing targeted attacks. We show that by carefully selecting the target, targeted attacks significantly outperform untargeted attacks in poisoning diffusion models and degrading the customization image quality. Extensive experiments validate the superiority of our method on two mainstream customization methods of diffusion models, compared to existing protections. To explain the surprising success of targeted attacks, we delve into the mechanism of attack-based protections and propose a hypothesis based on our observation, which enhances the comprehension of attack-based protections. To the best of our knowledge, we are the first to both reveal the vulnerability of diffusion models to targeted attacks and leverage targeted attacks to enhance protection against unauthorized diffusion customization.
@@ 2 Background
Diffusion Customization Diffusion customization (or diffusion personalization (Zhang et al., 2024)) refers to customizing the generated images by diffusion models according to the references from user inputs. These user inputs cover reference images (Meng et al., 2021), text description (Saharia et al., 2022), sketches (Zhang & Agrawala, 2023), styles (Ruiz et al., 2023; Zhang et al., 2023), and identities (Ruiz et al., 2023; Ye et al., 2023; Wang et al., 2024). Among these customization methods, LoRA+DreamBooth (Hu et al., 2021; Ruiz et al., 2023) and SDEdit (Meng et al., 2021) are two mainstream methods that cover most of the customization practice (See our brief survey in Appendix E). Hence, we follow the setup of existing works (Van Le et al., 2023; Liang et al., 2023; Xue et al., 2023) to focus on these two methods.
Adversarial Attacks & Protections against Unauthorized Diffusion Customization Adversarial attacks on Diffusion Models (Salman et al., 2023; Liang et al., 2023; Shan et al., 2023a; Liang & Wu, 2023; Van Le et al., 2023; Zhao et al., 2023; Xue et al., 2023; Liu et al., 2024b; a) have been used to protect private images against unauthorized diffusion customization. Specifically, these attacks add tiny adversarial perturbations to private images as protections. Diffusion models customized (fine-tuned (Hu et al., 2021; Ruiz et al., 2023) or referring to (Meng et al., 2021)) the protected images will produce images of degraded quality in customized generation. This will penalize the attempt to customize diffusion models on unauthorized private images.
It is noticeable that mainstream protections are based on untargeted attacks on diffusion models. Those untargeted objectives maximize the score-function error in Equation 2 by optimizing the image perturbation $\eta$ within the perturbation budget $\zeta$, which is formulated by Equation 3.
@@ 3.1 Attacking with Consistent Errors
$\left\{\begin{aligned} &\max_{\eta}\|p_{\theta}(y|x^{\prime})-\bm{e}_{true}\|_{2}^{2},\quad\quad\quad\quad&\text{(untargeted attack)}\\ &{\color[rgb]{1,.5,0}\min_{\eta}}\|p_{\theta}(y|x^{\prime})-{\color[rgb]{1,.5,0}\bm{e}_{target}}\|_{2}^{2},\quad\quad\quad\quad&\text{(targeted attack)}\end{aligned}\right$.
@@
$\left\{\begin{aligned} &\max_{\eta}\mathbb{E}_{t}\mathbb{E}_{z^{\prime}(t)|z^{\prime}(0)}\|s_{\theta}(z^{\prime}(t),t)-\nabla_{z^{\prime}(t)}\log p_{t}(z^{\prime}(t)|z^{\prime}(0))\|^{2}_{2},\quad\quad\quad\quad&\text{(untargeted attack)}\\ &{\color[rgb]{1,.5,0}\min_{\eta}}\mathbb{E}_{t}\mathbb{E}_{z^{\prime}(t)|z^{\prime}(0)}\|s_{\theta}(z^{\prime}(t),t)-{\color[rgb]{1,.5,0}\mathcal{T}}\|^{2}_{2},\quad\quad\quad\quad&\text{(targeted attack)}\end{aligned}\right$.
@@ 4 Experiments
We compare ACE/ACE+/ACE to existing protections against unauthorized diffusion customization in two main diffusion customization methods, LoRA+DreamBooth (Hu et al., 2021; Ruiz et al., 2023) and SDEdit (Meng et al., 2021). We also evaluate our protection on DreamBooth and Stable Diffusion 3 (Esser et al., 2024) (Appendix B.3). Our methods outperform baselines both in quantitative metrics and qualitative visualization (Section 4.2). We conduct user studies in an artist community, which validates the superiority of our methods (Section 4.3). Furthermore, we test the transferability (Section 4.4) and robustness to purification (Section 4.5) of ACE/ACE+. We investigate the computational cost of our methods and baselines (Appendix B.1). Our ablation studies focus on the effect of the objective function, target selection (Appendix B.5.1), different text prompts (Appendix B.5.2) and perturbation magnitudes (Appendix B.5.3) to the protection effectiveness.
@@ 4.1 Experimental Setups
Baselines We compare ACE/ACE+ with existing methods, including AdvDM (Liang et al., 2023), PhotoGuard (Salman et al., 2023), and ASPL (Anti-DreamBooth) (Van Le et al., 2023). Encoder Attack and Diffusion Attack in PhotoGuard (Salman et al., 2023) are denoted by PG and PG+, respectively. We detail our baseline selection in Appendix A
@@ 4.2 Overall Result
In addition to quantitative evaluation, we visualize the customization output images under different protection. Figure 1 showcases output images of LoRA and SDEdit under the protection of baselines and ACE/ACE+. We visualize the output images of ACE as an ablation study in Figure 11. Notably, all baseline methods only add chaotic patterns to some parts of the output image. Diffusion customization still succeeds in generating images with correct contents and styles. In contrast, ACE/ACE+ covers the whole output image with high contrast and overly busy patterns, making the image completely unusable. This advantage distinguishes ACE/ACE+ from existing protection against unauthorized diffusion customization. More visualization is given in Appendix F.
@@ 4.3 User Study
Protections against unauthorized diffusion customization aim at making the output image of unauthorized customization unusable in graphic art industries. The gold standard for evaluating protection is user studies by graphic artists. Hence, we conduct a survey in an artist community on the comparative effectiveness of two most competitive protections: ASPL (Van Le et al., 2023), and ACE. Basically, our survey asks participants to pick the image with the worst quality out of two customization images under the protection of ASPL, and ACE, respectively. The strongest protection should make the output image of the worst quality. We omit the details to Appendix E. The survey covers 1304 artists, showing that ACE produces the worst customization image among $55\%$ of the examples.
@@ 4.4 Transferability
Protections against unauthorized diffusion customization are based on white-box and model-dependent adversarial attacks. Hence, it is important to investigate the transferability of ACE/ACE+ over different diffusion models to validate their utility. We follow existing literature (Liang et al., 2023; Xue et al., 2023) to evaluate this transferability on three text-guided diffusion models: Stable Diffusion 1.4, Stable Diffusion 1.5, and Stable Diffusion 2.1. As mentioned in Section 2 and Section 4.1 and shown in Appendix E, we only consider these SD-family models because they cover most of unauthorized diffusion customization.
In this experiment, backbone means the model used to generate adversarial examples and victim means the model used in few-shot generation. We pick 100 images from CelebA-HQ, 20 in each of 5 groups, to generate adversarial examples on three backbone models and run SDEdit and LoRA with these adversarial examples on three victim models. The strength of SDEdit is set to 0.4 for visualizing the differences. Other experimental setups stay the same with Section 4.1. Table 2 shows that ACE/ACE+ can degrade diffusion customization even when the victim is different from the backbone. We omit the detailed explanation of the result to Appendix F.2.
@@ 5 Discussion: How can targeted attack beats untargeted attack?
While our methods empirically beat baselines in protecting images from unauthorized diffusion customization, it is non-trivial to explain our success. In this section, we discuss the process of fine-tuning diffusion models on protected images and provide one view of analysis on how these attack-based protections work and why ACE/ACE+/ACE perform distinguishably better than untargeted baselines on countering LoRA fine-tuning.
@@
2) $\mathcal{B}_{spl}$ is the score-function error of clean image $x$ in the fine-tuned diffusion model $\theta^{\prime}$.
$\left\{\begin{aligned} &\epsilon_{adv}:=\mathbb{E}_{t}\mathbb{E}_{z^{\prime}(t)|z^{\prime}(0)}\|s_{\theta}(z^{\prime}(t),t)-\nabla_{z^{\prime}(t)}\log p_{t}(z^{\prime}(t)|z^{\prime}(0))\|^{2}_{2}\\ &\mathcal{B}_{spl}:=\mathbb{E}_{t}\mathbb{E}_{z(t)|z(0)}\|s_{\theta^{\prime}}(z(t),t)-\nabla_{z(t)}\log p_{t}(z(t)|z(0))\|^{2}_{2}\end{aligned}\right$.
@@ Appendix A Implementation and Experimental Details
$\text{CLIP-SIM}(X,Y)={\color[rgb]{0,0,0}100\times}\cos(\mathcal{E}_{clip}(X),\mathcal{E}_{clip}(Y))$ (6)
@@
Baseline Selection Our works focuses on improving the protection effectiveness against unauthorized diffusion customization. Hence, we include baseline protection methods that 1) focus on effectiveness and 2) has open-sourced implementation or guidelines for reproduction. Also, to align with most baselines and make sure the comparison is fair, our evaluation is conducted on the same adversarial budget 3) under l2-constraint. Based on these three conditions, we select our baselines (Salman et al., 2023; Liang et al., 2023; Van Le et al., 2023). We notice that there are more existing works that contribute to the protection. However, some of them do not put their main contribution on improving effectiveness. For example, SDS (Xue et al., 2023) focuses on improving computational efficiency and MetaCloak (Liu et al., 2024b) focuses on robustness. Also, some existing works are lack of official open-sourced implementation (Liu et al., 2024a), while some others do not use l2-constraint (Shan et al., 2023a). Therefore, we do not include them in our baselines.
@@ B.2 Robustness to purification methods
Qualitatively, Table 3 also visualizes the output images of SDEdit and LoRA referring to the protected images processed by different purification methods. For both method, they still have strong performance under most of the cases except for Gaussian Noise($\sigma=8$) and JPEG compression (quality=20=20). However, in the exception cases, the defense has added visible degradation to the image, which also heavily affect both LoRA and SDEdit process. For example, LoRA learns to produce images comprised of small squares due to a hard compression of quality 20. And SDEdit produces images of visible Gaussian noise when adding Gaussian noise of $\sigma=8$. It’s noteworthy that both ACE and ACE+ seems to be strengthened after SR purification, which is an intriguing phenomena.
@@ B.3 Performance on Dreambooth
B.4 Performance on Stable Diffusion 3
To investigate whether ACE works on diffusion models based on Diffusion Transformers (DiTs) (Peebles & Xie, 2023) and Rectified Flow (Liu et al., 2022), we evaluate ACE on Stable Diffusion 3 (Esser et al., 2024), one state-of-the-art diffusion models for text-to-image generation. We follow our default setup on LoRA-DreamBooth, except for: 1) the image resolution is 1024 as the default of Stable Diffusion 3, 2) the adversarial budget is set to $8/255$, and 3) the training steps of LoRA are adapted to 2000 for training convergence. We train LoRA on both the clean and the protected subsets of CelebA-HQ. Then, we calculate the CLIP-IQA for both groups.
The CLIP-IQA of LoRA on the clean subset is $0.2650$, while that of LoRA on protected subset is $0.4963$. This validates that ACE is effective on Stable Diffusion 3. ACE is the first protection method that proves to work on Stable Diffusion 3 or other DiT-based diffusion models. Notably, ACE seems to work on Stable Diffusion 3 in a mechanism different from that on early versions of Stable Diffusion. Instead of degrading the quality of output images, ACE deviates the semantic of output images. For example, output images with LoRA trained on portraits of a person A will be portraits of other irrelevant persons. We visualize this phenomena in Figure 7. This means that one attack could perform in different ways on diffusion models with different architectures, which is a brand new finding about adversarial attacks on diffusion models. We leave its explanation to future work.
Figure 7: Comparison of output images from LoRA trained on clean images and protected images by ACE on Stable Diffusion 3. On Stable Diffusion 3, LoRA under ACE will deviate the semantics in the training images and generate irrelevant images, for example, generating portraits different from the portraits in the training images.
B.5 Ablation Studies
B.5.1 Targets and Objectives
We conduct ablation studies on the effect of targets and objectives. For targets, we use two natural images, one human-face photo from CelebA-HQ and one artwork from WikiArt, as alternative images for the target $\bm{\mathcal{T}}$. This is recommended by Van Le et al. (2023). For objectives, we use ASPL-T as the alternative objective. The experiment setup follows the LoRA setup in Section 4.1.
@@ Appendix C Limitations & Societal Impacts
In addition, we follow current works to use l2-distance to constrain our adversarial perturbations, which makes comparisons fair. This design, however, could be vulnerable to some new advanced purification methods, for example, IMPRESS (Cao et al., 2023) and GrIDPure (Zhao et al., 2024). We believe this could be overcome by rectifying the constraint and designing patterns that hide our protection deeper in the original image. However, this is out of the scope of this paper which focuses on optimizing the attack objective. We believe our stronger objective can contribute to real-world protection in combination with a potential robust constraint and pattern, where we leave to the future work. Also, our robustness experiments show that our method has robustness against general purification, keeping its essential utility. Hence, we leave the improvement of the robustness against advanced purification to the future work.
@@
Appendix D Key differences between our method and other targeted attack
ACE/ACE+ are not the first targeted attack on LDM. ASPL-T is a general targeted form of attacks (Van Le et al., 2023). Here, we compare the objective function of ACE and ASPL-T  [^13] The original paper does not provide the concrete objective function of ASPL-T so we refer to its implementation on GitHub: https://github.com/VinAIResearch/Anti-DreamBooth/blob/main/attacks/aspl.py:
Glaze (Shan et al., 2023a) is another well-known targeted attack. Two core designs of our method are both different from Glaze:
a) objective ACE minimizes the distance between diffusion predicted latent noise and a fixed latent. We have both the VAE $\mathcal{E}$ and the UNet $\theta$ (or Transformers in DiT architecture) involved. By contrast, Glaze minimizes the feature distance between the protected image and the target image in the VAE representation space. Only the VAE $\mathcal{E}$ is used in this process.
Glaze: $\mathop{min}_{x^{\prime}}\|z^{\prime}(0)-z_{\mathcal{T}}(0)\|^{2}_{2}$
$\text{where }z^{\prime}(0)=\mathcal{E}(x^{\prime}),z_{\mathcal{T}}(0)=\mathcal{E}(x_{\mathcal{T}}),\|x^{\prime}-x\|<p$
Here, $\|x^{\prime}-x\|$ is some distance metric and instantiated as LPIPS (Zhang et al., 2018) in the implementation.
b) target ACE exploits the latent of a chaotic-pattern image $x_{\mathcal{T}}$ as the target. Glaze uses the style-transfered version of the original image as the target image.
Glaze: $x_{\mathcal{T}}=\Omega(x,T)$ (9)
[…]
Appendix E Toy Survey on Diffusion Customization